# Handling Domain Shifts for Anomalous Sound Detection: A Review of DCASE-Related Work

### v2

@@ Abstract
Index Terms:
anomalous sound detection, domain shift, domain adaptation, domain generalization, machine condition monitoring
@@ 1 Introduction
# recordings (per section) # machine types # sections (per machine type) source domain target domain Name Purpose total dev. set eval. set total dev. set eval. set split normal anomalous normal anomalous supplemental DCASE2021 [1] DA $7$ 7 7 6 3 3 train $1000$ 00 $3$ 00 00 DCASE2021 [1] DA $7$ 7 7 6 3 3 test $100$ $100$ $100$ $100$ 00 DCASE2022 [2] DG $7$ 7 7 6 3 3 train $990$ 00 $10$ 00 00 DCASE2022 [2] DG $7$ 7 7 6 3 3 test $50$ $50$ $50$ $50$ 00 DCASE2023 [3] DG $14$ 7 7 1 1 1 train $990$ 00 $10$ 00 00 DCASE2023 [3] DG $14$ 7 7 1 1 1 test $50$ $50$ $50$ $50$ 00 DCASE2024 [4] DG $16$ 7 9 1 1 1 train $990$ 00 $10$ 00 00 DCASE2024 [4] DG $16$ 7 9 1 1 1 test $50$ $50$ $50$ $50$ 00 DCASE2025 [5] DG $14$ 7 7 1 1 1 train $990$ 00 $10$ 00 $100$ DCASE2025 [5] DG $14$ 7 7 1 1 1 test $50$ $50$ $50$ $50$ 00
@@ 2 Domain Shifts
The goal of ASD is to determine, based on a sound recording of a phenomenon, whether that phenomenon is normal or anomalous. Its normal/anomalous characteristic is an intrinsic property of that phenomenon: as such, the determination by the ASD system should ideally be independent of the conditions in which the sound was recorded, as long as the change in conditions does not alter the observability and nature of the phenomenon’s normal/anomalous characteristic. Such changes include the use of microphones of different types or in different locations, the presence or absence of other sound sources, or modifications in certain properties of the monitored sound sources themselves, e.g., the use of different settings of the monitored machines.
@@ 3.1 Datasets
[…]
@@ 4 Domain Generalization
Adapting models for each unknown domain with the possible need of re-training models, fine-tuning hyperparameters, or even replacing system components is very impractical, as it is computationally costly and may require expert knowledge. It is thus much more desirable to strive for domain generalization (DG) [40, 41], with the goal of obtaining a model that performs well in the source domain and generalizes well to unseen target domains by only providing a few samples to define how normal sounds belonging to these domains should sound like. Since DG is literally a generalization of DA to arbitrary domains rather than a specific target domain, it is more difficult to obtain such a model.
@@ 4.2.4 Anomaly Score Calculation
Last but not least, DG capabilities can also be improved by modifying the anomaly score computation. This approach has the advantage of not requiring expensive re-training of the neural networks that serve as the basis of state-of-the-art ASD systems. To this end, [32] uses an autoencoder and calculates the Mahalanobis distance between input data and reconstruction using domain-specific covariance matrices. For discriminative ASD systems, it was shown that simple nearest-neighbor-based anomaly scores lead to better results than estimating domain-specific distributions [33]. Further improvements can be obtained by normalizing the anomaly scores to reduce the domain mismatch. Examples are normalizing the anomaly scores based on local densities of normal reference samples [34] or by a domain-wise standardization of the anomaly score distributions [35].
@@ 6 Conclusion
In this work, we reviewed recent work on how to handle domain shifts for ASD tasks related to DCASE. We first defined domain shifts in the context of ASD, then discussed the topics of domain adaptation (DA) and domain generalization (DG) by motivating and defining these terms, presenting relevant datasets, and collecting different works related to these topics. In particular, DG approaches were grouped into domain specialization, domain-invariant representations, feature disentanglement, and anomaly score calculation. Furthermore, continuous and continual DA as well as explainable DG were identified as possible future research directions. For future work, we plan to compare all presented techniques in extensive experimental evaluations to identify the best working approaches and possible synergies.

### v3

@@ 1 Introduction
# recordings (per section) # machine types # sections (per machine type) source domain target domain Name Purpose total dev. set eval. set total dev. set eval. set split normal anomalous normal anomalous supplemental DCASE2021 [1] DA 7 7 7 6 3 3 train $1000$ 0 33 0 0 test $100$ $100$ $100$ $100$ 0 DCASE2022 [2] DG 7 7 7 6 3 3 train $990$ 0 $10$ 0 0 test $50$ $50$ $50$ $50$ 0 DCASE2023 [3] DG 14 7 7 1 1 1 train $990$ 0 $10$ 0 0 test $50$ $50$ $50$ $50$ 0 DCASE2024 [4] DG 16 7 9 1 1 1 train $990$ 0 $10$ 0 0 test $50$ $50$ $50$ $50$ 0 DCASE2025 [5] DG 14 7 7 1 1 1 train $990$ 0 $10$ 0 $100$ test $50$ $50$ $50$ $50$ 0
@@ 2 Domain Shifts
The goal of ASD is to determine, based on a sound recording of a phenomenon, whether that phenomenon is normal or anomalous. In the context of the DCASE Challenges, the objective is to differentiate between normal and anomalous sounds produced by a known type of machine. The normal/anomalous characteristic of a phenomenon is an intrinsic property of that phenomenon: as such, the determination by the ASD system should ideally be independent of the conditions in which the sound was recorded, as long as the change in conditions does not alter the observability and nature of the phenomenon’s normal/anomalous characteristic. Such changes include the use of microphones of different types or in different locations, the presence or absence of other sound sources, or modifications in certain properties of the monitored sound sources themselves, e.g., the use of different settings of the monitored machines.
@@ 3.1 Datasets
A dataset focusing on handling domain shifts for ASD through DA is the DCASE2021 ASD dataset [1]. This dataset contains $10\text{\,}\mathrm{s}$ recordings of five different machine types from MIMII DUE [13] and two additional machine types from ToyADMOS2 [14] that are combined with background noise from real factories. The dataset is divided into a development set and an evaluation set, which both contain three sections for each machine type. These sections are specific partitions of the dataset for calculating the performance and may contain recordings from multiple machines of the same type. Both the development and evaluation sets consist of a training split and a test split. The training splits contain only normal data, of which 1,0001{,}000 samples belong to the source domain and only $3$ samples belong to the target domain, resulting in a very imbalanced dataset in terms of domain. For each training file, additional attribute information about machine settings or the acoustic environment is provided, and it can be utilized for training ASD systems. The test splits contain $200$ samples for each domain, of which one half are normal and the other half are anomalous. For each test file, it is known whether these files belong to the source or the target domain, but it is unknown whether they are normal or anomalous. Summarized statistics of this dataset can be found in the top part of Table 1.
@@ 4 Domain Generalization
Adapting models for each new domain with the possible need of re-training models, fine-tuning hyperparameters, or even replacing system components is very impractical, as it is computationally costly and may require expert knowledge. It is thus much more desirable to strive for domain generalization (DG) [40, 41], with the goal of obtaining a model that performs well in the source domain and generalizes well to unseen target domains by only providing a few samples to define how normal sounds belonging to these domains should sound like. Since DG is literally a generalization of DA to arbitrary domains rather than a specific target domain, it is more difficult to obtain such a model. To highlight the practical differences between DA and DG, the reader is reminded that for DA, models can be fully adapted to a specific target domain, even if performance in the source domain is weak after adaptation. In contrast, DG requires strong performance in the source and target domain, using even the same decision threshold for the anomaly scores in both domains.
@@ 4.2.4 Anomaly Score Calculation
Last but not least, DG capabilities can also be improved by modifying the anomaly score computation. This approach has the advantage of not requiring expensive re-training of the neural networks that serve as the basis of state-of-the-art ASD systems. To this end, [32] uses an autoencoder and calculates the Mahalanobis distance between input data and reconstruction using domain-specific covariance matrices. For discriminative ASD systems, it was shown that simple nearest-neighbor-based anomaly scores lead to better results than estimating domain-specific distributions [33]. Further improvements can be obtained by normalizing the anomaly scores to reduce the domain mismatch. Examples are normalizing the anomaly scores based on local densities of normal reference samples [55, 34] or by a domain-wise standardization of the anomaly score distributions [35].
@@ 6 Conclusion
In this work, we reviewed recent work on how to handle domain shifts for ASD tasks related to DCASE. We first defined domain shifts in the context of ASD, then discussed the topics of domain adaptation (DA) and domain generalization (DG) by motivating and defining these terms, presenting relevant datasets, and collecting different works related to these topics. In particular, DG approaches were grouped into domain specialization, domain-invariant representations, feature disentanglement, and anomaly score calculation. Furthermore, continuous and continual DA as well as explainable DG were identified as possible future research directions. For future work, we plan to compare all presented techniques with ASDKit [64] through extensive experiments to identify the best approaches and potential synergies.